How much work was required to bring the 1000-kg roller coaster from point P to rest at point Q at the top of the 50-m peak?
1
Expert's answer
2013-02-27T04:48:16-0500
The work is equal to the change of potential energy V ofthe coaster.

The value of V at the height h is equal to V(h) = m*g*h,
where g=9.8 m/s^2. Thus,
A = V(h) - V(0)
= m * g * h - 0 =
= m * g * h =
= 1000 * 9.8 * 50
= 490000 J.